# Imaging the initial condition of heavy-ion collisions and nuclear structure across the nuclide chart

### v2

@@ Abstract
High-energy nuclear collisions encompass three key stages: the structure of the colliding nuclei, informed by low-energy nuclear physics, the initial condition, leading to the formation of quark-gluon plasma (QGP), and the hydrodynamic expansion and hadronization of the QGP, leading to final-state hadron distributions that are observed experimentally. Recent advances in both experimental and theoretical methods have ushered in a precision era of heavy-ion collisions, enabling increasingly accurate understanding of these stages. However, most approaches involve simultaneously determining both QGP properties and initial conditions from a single collision system, creating complexity due to the coupled contributions of these stages to the final-state observables.
To avoid this, we propose leveraging established knowledge of low-energy nuclear structure and hydrodynamic observables to independently constrain the QGP’s initial condition. By conducting comparative studies of collisions involving isobar-like nuclei – species with similar mass numbers but different ground-state geometries – we can disentangle the initial condition’s impacts from the QGP properties. This approach not only refines our understanding of the initial stages of the collisions but also turns high-energy nuclear experiments into a precision tool for imaging nuclear structure, offering insights that complement traditional low-energy approaches.
Opportunities for carrying out such comparative experiments at the Large Hadron Collider (LHC) and other facilities could significantly advance both high-energy and low-energy nuclear physics. Additionally, this approach holds potential implications for the future Electron-Ion Collider (EIC). While the possibilities are extensive, we focuses on selected proposals that could benefit both the high-energy and low-energy nuclear physics communities. Originally prepared as input for the long-range plan of U.S. nuclear physics, this white paper reflects the status as of September 2022, with a brief update on developments since then.
@@ I Introduction: from nuclear structure to heavy-ion collisions
A major research goal in high-energy nuclear physics is the characterization of the QGP in terms of medium properties, such as specific shear and bulk viscosities, $\eta/s$ and $\zeta/s$, or the jet quenching transport parameter, $\hat{q}$ Busza et al. (2018). The precision achievable in this characterization, e.g., in state-of-the-art Bayesian analyses Bernhard et al. (2016); Everett et al. (2021); Nijs et al. (2021); Xie et al. (2022); Ehlers et al. (2022), is impacted by our uncertain knowledge of the mechanism of energy deposition in the interaction of two nuclei. Assessing the role of the nuclear structure input will, therefore, reduce this uncertainty, permitting global analyses of data to infer cleaner information about the collision dynamics, and in turn the knowledge of the QGP initial condition Giacalone et al. (2021a). Conversely, a major direction of research in nuclear structure theory focuses on the emergence of nuclear properties from fundamental theory Hergert (2020). Such ab initio approaches aim at describing strongly-correlated nuclear systems from approximate (yet systematically improvable) solutions of the Schrödinger equation with nucleon-nucleon and three-nucleon interactions constructed in an effective theory of low-energy QCD. These efforts find a natural application in the phenomenology of multi-particle correlations in high-energy nuclear collisions. Once the response of the QGP initial condition to nuclear structure is established, one could use measurements in heavy-ion collisions to test the results of ab initio approaches in a way that is complementary to low-energy experiments. The systematic use of ab initio results as an input for the model building of nuclear collisions will then permit us to assess, in particular, the consistency of nuclear phenomena across energy scales.
@@ IV Brief summary of developments since 2022
One notable development is the progress made in extracting quantitative information about the structure of nuclei from collider data. Using the Bayesian analysis framework, a recent study demonstrated the possibility of determining the skin of 208Pb from LHC data Giacalone et al. (2023), yielding values that align well with low-energy determinations Adhikari et al. (2021); Hu et al. (2022). Additionally, quantitative insights into nuclear deformations were obtained through comparative measurements of 238U+238U and 197Au+197Au collisions conducted by the STAR Collaboration STA (2024). When combined with high-precision hydrodynamic calculations and the relatively well-understood, modest quadrupole deformation of 197Au Bally et al. (2023); Ryssens et al. (2023), these measurements enabled the quantitative extraction of the intrinsic surface deformation parameters, $\beta_{2\rm U}$ and $\gamma_{\rm U}$, of the 238U nucleus. The results for $\beta_{2\rm U}$ align with low-energy nuclear physics results, while the non-zero value of $\gamma_{\rm U}$ provides evidence for a shape that breaks axial symmetry, shedding light on an aspect of this nucleus that is poorly known in low-energy approaches.
@@ Acknowledgments
We thank the participants of the EMMI Rapid Reaction Task Force "Nuclear physics confronts relativistic collisions of isobars" (https://indico.gsi.de/event/14430/) for valuable input.

### v4

@@ Abstract
High-energy nuclear collisions encompass three key stages: the structure of the colliding nuclei, informed by low-energy nuclear physics, the initial condition, leading to the formation of quark-gluon plasma (QGP), and the hydrodynamic expansion and hadronization of the QGP, leading to final-state hadron distributions that are observed experimentally. Recent advances in both experimental and theoretical methods have ushered in a precision era of heavy-ion collisions, enabling an increasingly accurate understanding of these stages. However, most approaches involve simultaneously determining both QGP properties and initial conditions from a single collision system, creating complexity due to the coupled contributions of these stages to the final-state observables.
To avoid this, we propose leveraging established knowledge of low-energy nuclear structures and hydrodynamic observables to independently constrain the QGP’s initial condition. By conducting comparative studies of collisions involving isobar-like nuclei – species with similar mass numbers but different ground-state geometries – we can disentangle the initial condition’s impacts from the QGP properties. This approach not only refines our understanding of the initial stages of the collisions but also turns high-energy nuclear experiments into a precision tool for imaging nuclear structures, offering insights that complement traditional low-energy approaches.
Opportunities for carrying out such comparative experiments at the Large Hadron Collider (LHC) and other facilities could significantly advance both high-energy and low-energy nuclear physics. Additionally, this approach has implications for the future Electron-Ion Collider (EIC). While the possibilities are extensive, we focus on selected proposals that could benefit both the high-energy and low-energy nuclear physics communities. Originally prepared as input for the long-range plan of U.S. nuclear physics, this white paper reflects the status as of September 2022, with a brief update on developments since then.
@@ I Introduction: from nuclear structure to heavy-ion collisions
A major research goal in high-energy nuclear physics is the characterization of the QGP in terms of medium properties, such as specific shear and bulk viscosities, $\eta/s$ and $\zeta/s$, or the jet quenching transport parameter, $\hat{q}$ Busza et al. (2018). The precision achievable in this characterization, e.g., in state-of-the-art Bayesian analyses Bernhard et al. (2016); Everett et al. (2021); Nijs et al. (2021); Xie et al. (2023); Ehlers (2023), is impacted by our uncertain knowledge of the mechanism of energy deposition in the interaction of two nuclei. Assessing the role of the nuclear structure input will, therefore, reduce this uncertainty, permitting global analyses of data to infer cleaner information about the collision dynamics, and in turn the knowledge of the QGP initial condition Giacalone et al. (2021a). Conversely, a major direction of research in nuclear structure theory focuses on the emergence of nuclear properties from fundamental theory Hergert (2020). Such ab initio approaches aim at describing strongly-correlated nuclear systems from approximate (yet systematically improvable) solutions of the Schrödinger equation with nucleon-nucleon and three-nucleon interactions constructed in an effective theory of low-energy QCD. These efforts find a natural application in the phenomenology of multi-particle correlations in high-energy nuclear collisions. Once the response of the QGP initial condition to nuclear structure is established, one could use measurements in heavy-ion collisions to test the results of ab initio approaches in a way that is complementary to low-energy experiments. The systematic use of ab initio results as an input for the model building of nuclear collisions will then permit us to assess, in particular, the consistency of nuclear phenomena across energy scales.
@@ IV Brief summary of developments since 2022
One notable development is the progress made in extracting quantitative information about the structure of nuclei from collider data. Using the Bayesian analysis framework, a recent study demonstrated the possibility of determining the skin of 208Pb from LHC data Giacalone et al. (2023), yielding values that align well with low-energy determinations Adhikari et al. (2021); Hu et al. (2022). A 3+1D simulations of isobar collisions have also revealed the sensitivity to the neutron skin of Ru and Zr nuclei in observables that compare baryon to electric charge stopping Pihan et al. (2024). Additionally, quantitative insights into nuclear deformations were obtained through comparative measurements of 238U+238U and 197Au+197Au collisions conducted by the STAR Collaboration STAR Collaboration (2024). When combined with high-precision hydrodynamic calculations and the relatively well-understood, modest quadrupole deformation of 197Au Bally et al. (2023); Ryssens et al. (2023), these measurements enabled the quantitative extraction of the intrinsic surface deformation parameters, $\beta_{2\rm U}$ and $\gamma_{\rm U}$, of the 238U nucleus. The results for $\beta_{2\rm U}$ align with low-energy nuclear physics results, while the non-zero value of $\gamma_{\rm U}$ provides evidence for a shape that breaks axial symmetry, shedding light on an aspect of this nucleus that is poorly known in low-energy approaches.
@@ Acknowledgments
We thank the participants of the EMMI Rapid Reaction Task Force "Nuclear physics confronts relativistic collisions of isobars" (https://indico.gsi.de/event/14430/) for valuable input. This work is supported by U. S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award or Contract No. DE-SC002418 (JDB), DE-SC0024602 (SH, JJ), DE-SC0004286 (UH), DE-FG02-10ER41666 (CL, WL), DE-SC0013365, DE-SC0024586 and DE-SC0023175 (DL), DE-SC0011088 (YL), DE-AC02-05CH11231 (MP), DE-FG02-89ER40531 (AT), DE-SC0012704 (BS), DE-SC0021969 and DE-SC0024232 (CS), DE-SC0023861 (JN), DE-FG02-07ER41521 (ZX); by National Science Foundation under grant number OAC-2103680 (JN); by European Union (ERC, Initial Conditions), VILLUM FONDEN with grant no. 00025462, and Danmarks Frie Forskningsfond (YZ); and by FAPESP projects 2017/05685-2, 2018/24720-6, and 2021/08465-9, project INCT-FNA Proc. No. 464898/2014-5, and CAPES - Finance Code 001 (ML).